Differential Equations

Solve this Initial Value Problem
X۔ 5= (0)ی 2 = (0) لارج (+3) = 9+ و +
Answer #1

Given y +4y + 4y = (3+x) = 24 ز 40)=2; y co-5 2. (b + 4.0+4) y = (3+2) €27; fez o²+ 4044 ; Q6A) = (3+1)ě?? Auxilary Equatio3ta $ è 2x [ € 3+x)] +*(1+2 A C (-2:30 - -) (3+)] = [(3+4)=40,1)] ē 27 = [x-8 16 Complete solution g=yc typ (citc27) e2x + -25 = 5 3 C2+1 -2 [ air 5 - Catto -2(2) q=catto (2=9-to (2= 143 16 complete solution 288 28f 1437 16 5 32 -24 =e 69 +2437*2)+2x